OpenStudy (vannah):

CAN ANYONE GIVE ME AN EXPLANATION TO THIS PROBLEM AND HOW THEY GOT THE ANSWER 2X-3-15X+6

OpenStudy (anonymous):

2x-3-15x+6 ?

OpenStudy (vannah):

Yeah that

OpenStudy (anonymous):

add like term 2x-15x -3+6

OpenStudy (anonymous):

2-15=-13 -3+6=3 -13x +3

OpenStudy (vannah):

okay thankyou